OpenStudy (anonymous):

Please help! I'll give you a medal<3 1. Which ordered pair is NOT a solution of y = –12x – 3? A. (-3,33) B. (8, -98) C. (-9, 105) D) (3, -39) (Reputing this question, to make sure!)

OpenStudy (anonymous):

A) -12(-3)=36-3=33 coordinates would be (-3,33) therefore this coordinate is correct. B) -12(8)=-96-3=99 coordinates would be (8,-99) therefore the answer to this question is B. C) -12(-9)=108-3=105 coordinates would be (-9,105) therefore this coordinate is correct. D) -12(3)=-36-3=-39 coordinates would be (3,-39) therefore this coordinate is correct.

OpenStudy (anonymous):

Thank you!<3

OpenStudy (anonymous):

no problem, i made a mistake there on B "-12(8)=-96-3=99" should be -99 but either way the answer is still B xD